# On the speed of coming down from infinity for (sub)critical branching processes with pairwise interactions

### v1

@@ 3.1 Integral representation of BPI processes
The proof is a simple adaptation of the argument used in the [13, Proof of Theorem 5.5]. Let $Y_{t}=Z^{(z_{1})}_{t}-Z^{(z_{2})}_{t}$, for $t\geq 0$. By Proposition 1,
$Y_{t}=z_{1}-z_{2}+\int_{0}^{t}\int_{\mathcal{X}}(H(Z^{(z_{1})}_{s-},u,y,x)-H(Z^{(z_{2})}_{s-},u,y,x))N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x).$ (20)
Let $(a_{m})_{m\geq 0}$ be a decreasing sequence of positive real number such that $a_{0}=1$, $a_{m}>0$ for $m\in\mathbb{N}$, and $\lim_{m\rightarrow\infty}a_{m}=0$. For each $m\in\mathbb{N}$, let $g_{m}:\mathbb{R}\rightarrow\mathbb{R}_{+}$ be a non-negative continuous function with support on $(a_{m},a_{m-1})$ and satisfying $\int_{a_{m}}^{a_{m-1}}g_{m}(z){\rm d}z=1$. For each $m\in\mathbb{N}$, we define the non-negative and twice continuously differentiable function
$\phi_{m}(z)=\int_{0}^{z}\int_{0}^{y}g_{m}(x){\rm d}x{\rm d}y,\qquad z\in\mathbb{R}.$
@@
$G_{s}(\phi_{m},Y)\coloneqq\phi_{m}(Y_{s-}+H(Z^{(z_{1})}_{s-},u,y,x)-H(Z^{(z_{2})}_{s-},u,y,x))-\phi_{m}(Y_{s-})=0,\quad s\geq 0.$ (21)
Note that $Y_{0}=z_{1}-z_{2}\leq 0$ and $\phi_{m}(Y_{0})=0$. For $n\in\mathbb{N}$, define $\tau_{n}=\inf\{t\geq 0:Z^{(z_{1})}_{t}\geq n\,\,\text{or}\,\,Z^{(z_{2})}_{t}\geq n\}$. Then, by (20), Itô’s formula ([18, Chapter 2, Theorem 5.1]) and (21),
@@
$=\int_{0}^{t\wedge\tau_{n}}\int_{\mathcal{X}}G_{s}(\phi_{m},Y)\mathbf{1}_{\{Y_{s-}>0\}}{\rm d}sq({\rm d}u,{\rm d}y,{\rm d}x)+M_{t}^{(n,m)},$ (22)
where $(M_{t}^{(n,m)})_{t\geq 0}$ is martingale. Note that $|\phi_{m}^{\prime}(z)|\leq 1$, for $z\in\mathbb{R}$, which implies that $\phi_{m}$ is (globally) Lipschitz. Since we have assumed that $\sum_{i\geq 1}i\pi_{i}<\infty$ and $\sum_{i\geq 1}ib_{i}<\infty$, it follows from (17) and (19), for any $s\leq\tau_{n}$,
$\int_{\mathcal{X}}G_{s}(\phi_{m},Y)\mathbf{1}_{\{Y_{s-}>0\}}q({\rm d}u,{\rm d}y,{\rm d}x)\leq\Big{(}d+(2n-1)c+\sum_{i\geq 1}i(\pi_{i}+(2n-1)b_{i})\Big{)}Y_{s-}^{+}.$ (23)
Then, from (3.1), (23) and the dominated convergence theorem (letting $m\rightarrow\infty)$, we deduce that
$\mathbb{E}[Y_{t\wedge\tau_{n}}^{+}]\leq\Big{(}d+(2n-1)c+\sum_{i\geq 1}i(\pi_{i}+(2n-1)b_{i})\Big{)}\int_{0}^{t}\mathbb{E}[Y_{s\wedge\tau_{n}}^{+}]{\rm d}s.$
Since $\int_{0}^{\varepsilon}u^{-1}{\rm d}u=\infty$, for $\varepsilon>0$, we conclude that $\mathbb{E}[Y_{t\wedge\tau_{n}}^{+}]=0$, for all $t\geq 0$. On the other hand, since $(Z^{(z_{1})}_{t})_{t\geq 0}$ and $(Z^{(z_{2})}_{t})_{t\geq 0}$ do not explode in finite time and have càdlàg paths, we deduce that $\tau_{n}\rightarrow\infty$, as $n\rightarrow\infty$. Therefore, our claim follows. ∎
Lemma 1.
Suppose that $\sum_{i\geq 1}i\pi_{i}<\infty$ and $\sum_{i\geq 1}ib_{i}<\infty$. Then, the BPI-process is stochastically monotone, that is, for all $z_{1},z_{2}\in\mathbb{N}_{0}$ such that $z_{1}\leq z_{2}$, $\mathbb{P}(Z_{t}^{(z_{1})}\geq x)\leq\mathbb{P}(Z_{t}^{(z_{2})}\geq x)$, for all $t\geq 0$ and $x\in\mathbb{R}$.
@@
It follows from Proposition 2. ∎
3.2 Proof of Theorem 1
In this section we prove Theorem 1 using the framework developed in [3, Section 4] (with $F(z)\coloneqq\log(1+z)$, for $z\in(-1,\infty)$). Recall that $\mathcal{X}\coloneqq\{1,2\}\times(\mathbb{N}\cup\Delta)\times\mathbb{N}_{-1}$. The next lemma verifies part of [3, Assumptions 4.1 and 4.3]. We have removed the subscript $F$ from our notation and will use $\varphi$ to denote the function $b_{F}$ in [3].
Lemma 2.
@@ 3.2 Proof of Theorem 1
    Define the function $h(z)\coloneqq\int_{\mathcal{X}}\log\Big{(}1+\frac{H(z,u,y,x)}{z+1}\Big{)}q({\rm d}u,{\rm d}y,{\rm d}x)$, for $z\in\mathbb{N}_{0}$. Then, $h$ can be extended to $[0,\infty)$ such that is locally bounded on $[0,\infty)$ and locally Lipschitz on $(0,\infty)$.
@@
    Define the function $\psi(z)\coloneqq e^{-z}\varphi(e^{z}-1)$, for $z\in[0,\infty)$. Then, $\psi$ is $(0,\alpha)$ nonexpansive on $(1,\infty)$ for some $\kappa\geq 0$, i.e. for all $z_{1}>z_{2}>0$, $\psi(z_{1})\leq\psi(z_{2})+\kappa$.
@@
$\ln(1+x)\leq x.$ (24)
First, we prove (i). If $z=0$, then $H(0,u,y,x)=0$, for all $u=1,2$, $y\in\mathbb{N}\cup\Delta$ and $x\in\mathbb{N}_{-1}$ and thus, (i) holds. For $z\in\mathbb{N}$, (i) follows from (17), (19), (24) and our assumptions $\mathbf{m}_{\rm int}<0$ and $\mathbf{m}_{\rm br}<\infty$, that is,
@@
$\leq d+\sum_{i\geq 1}i\pi_{i}+zc+z\sum_{i\geq 1}ib_{i}<\infty.$ (25)
Next, we prove (ii). Note that by (17) and (19),
$h(z)$ $=z\int_{\mathbb{N}_{-1}}\log\Big{(}1+\frac{x}{z+1}\Big{)}\nu_{\rm br}({\rm d}x)+\frac{z(z-1)}{2}\int_{\mathbb{N}_{-1}}\log\Big{(}1+\frac{x}{z+1}\Big{)}\nu_{\rm int}({\rm d}x)=h_{\rm br}(z)+h_{\rm int}(z),$ (26)
@@
$h_{\rm int}(z):=-z(z-1)c\log\Big{(}1+\frac{1}{z}\Big{)}+z(z-1)\sum_{i\geq 1}b_{i}\log\Big{(}1+\frac{i}{z+1}\Big{)}.$ (28)
Clearly, $h$ (and thus, $h_{\rm br}$ and $h_{\rm int}$) can be extended to $[0,\infty)$. In particular, (3.2) remains valid and it follows that $h$ is locally bounded on $[0,\infty)$. On the other hand, note that, for $0<x\leq y$, (24) implies that
$|\log(x)-\log(y)|=\log\left(\frac{y}{x}\right)=\log\left(1+\left(\frac{y}{x}-1\right)\right)\leq\frac{y-x}{x}.$ (29)
Then, by (29) and our assumptions that $\mathbf{m}_{\rm int}<0$ and $\mathbf{m}_{\rm br}<\infty$, it is not difficult to deduce that $h$ is locally Lipschitz on $(0,\infty)$, which concludes with the proof of (ii).
We now prove (iii). By (24) and since $\mathbf{m}_{\rm br}<\infty$, $h_{\rm br}$ is a bounded function on $[0,\infty)$. Note that for $x,y>0$, $\log(x)\leq\log(y)+(x-y)/y$. Then, for $z\in[0,\infty)$, $h_{\rm int}(z)\leq z\mathbf{m}_{\rm int}$. In particular, since $\mathbf{m}_{\rm int}<0$, $h_{\rm int}(z)\rightarrow-\infty$, as $z\rightarrow\infty$, which implies (iii).
@@
$|1/x-1/y|=\frac{|x-y|}{xy}.$ (30)
Then, by the triangle inequality and (30), we have that, for $z\in(1,\infty)$,
$\Big{|}-z(z-1)c\log(1+1/z)+cz\Big{|}$ $\leq z(z-1)\Big{|}c\int_{0}^{-1}\Big{(}\frac{1}{z+1+y}-\frac{1}{z-1}\Big{)}{\rm d}y\Big{|}\leq c.$ (31)
Recall that $h_{\rm br}$ is bounded on $[0,\infty)$ and thus, (31) implies that $\hat{h}$ is a bounded function on $(1,\infty)$. Moreover, $z\mapsto-cz+h_{\rm int}(z)+z(z-1)c\log(1+1/z)$ is a decreasing function on $(1,\infty)$. Then, (iv) follows with $\kappa=2||\hat{h}||_{\infty}$. ∎
Let $z_{0}>1$ be such that (iii) in Lemma 2 is satisfied. Recall the definition of the function $\varphi$ in Lemma 2 (iii). By Lemma 2 (i)-(ii), $\varphi$ is well-defined and locally Lipschitz on $D\coloneqq(z_{0},\infty)$. We introduce the flow $\phi$ associated to $\varphi$ and defined for $z\in D$ as the unique solution of
$\quad\frac{\partial}{\partial t}\phi(z,t)=\varphi(\phi(z,t)),\qquad\textrm{with}\quad\phi(z,0)=z,$ (32)
@@
$v_{t}\coloneqq\phi(\infty,t)=\lim_{z\rightarrow\infty}\phi(z,t),$ (33)
for any $t\in[0,T(\infty))$. By Lemma 2 (iii), for $z\in D$, $\varphi(z)<0$ and for any $t<T(z)$,
$\int_{z}^{\phi(z,t)}\frac{{\rm d}x}{\varphi(x)}=t.$
Lemma 3.
@@
$-\int_{z}^{\infty}\frac{{\rm d}x}{\varphi(x)}<\infty.$
@@
The first claim follows from the definition of $\varphi$, (26) and the dominated convergence theorem. The second claim follows clearly from the first claim. ∎
By Lemma 3, we have that $v_{t}$ satisfies
$-\int_{v_{t}}^{\infty}\frac{{\rm d}x}{\varphi(x)}=t,$ (34)
@@
$v_{t}=\inf\Big{\{}z\in D:\int_{z}^{\infty}\frac{{\rm d}x}{-\varphi(x)}<t\Big{\}}<\infty,$ (35)
for any $t\in(0,T(\infty))$. Following [3, Section4], one says that the dynamical system instantaneously comes-down from infinity. The function $t\in[0,T(\infty))\rightarrow v_{t}\in\overline{\mathbb{R}}$ is continuous when $\overline{\mathbb{R}}$ is endowed with the distance $\overline{d}(x,y)=|e^{-x}-e^{-y}|$, for $x,y\in\overline{\mathbb{R}}$.
Lemma 4.
@@
By (33), $v_{t}\in D$ and thus $v_{t}>0$, for $t\in[0,T(\infty))$. By Lemma 2 (iii), $-\varphi(\cdot)$ is strictly positive on $D$. Then, Lemma 3 implies that $z\mapsto g(z)=-\int_{z}^{\infty}\frac{{\rm d}x}{-\varphi(x)}$ maps $D=(z_{0},\infty)$ bijectively to $(0,-\int_{z_{0}}^{\infty}\frac{{\rm d}x}{-\varphi(x)})$. Moreover, $g$ is decreasing on $(z_{0},\infty)$. Then, it should be clear that $\lim_{t\downarrow 0}v_{t}=\infty$, which proves (i). (ii) follows by (34) and the fundamental theorem of calculus; recall that for $z\in D$, $\varphi(z)<0$. From (34), we see that $tv_{t}=v_{t}\int_{v_{t}}^{\infty}\frac{{\rm d}x}{-\varphi(x)}$, for $t\in(0,T(\infty))$. Then, L’Hôpital’s rule and Lemma 3 imply (iii), that is,
$\lim_{t\downarrow 0}-\mathbf{m}_{\rm int}tv_{t}=-\mathbf{m}_{\rm int}\lim_{z\rightarrow\infty}z\int_{z}^{\infty}\frac{{\rm d}x}{-\varphi(x)}=-\mathbf{m}_{\rm int}\lim_{z\rightarrow\infty}\frac{\frac{1}{\varphi(z)}}{-\frac{1}{z^{2}}}=1.$ (36)
@@
The following lemma shows that condition [3, (21) in Assumption 4.3] is satisfied.
Lemma 5.
@@
$V(z)\coloneqq\int_{\mathcal{X}}\Big{(}\log\Big{(}1+\frac{H(z,u,y,x)}{z+1}\Big{)}\Big{)}^{2}q({\rm d}u,{\rm d}y,{\rm d}x),$ (37)
@@
$\int_{0}^{t}\widehat{V}_{\varepsilon}(z_{0},s){\rm d}s<\infty,$ (38)
@@
$V(z)=z\int_{\mathbb{N}_{-1}}\Big{(}\log\Big{(}1+\frac{x}{z+1}\Big{)}\Big{)}^{2}\nu_{\rm br}({\rm d}x)+\frac{z(z-1)}{2}\int_{\mathbb{N}_{-1}}\Big{(}\log\Big{(}1+\frac{x}{z+1}\Big{)}\Big{)}^{2}\nu_{\rm int}({\rm d}x).$ (39)
On the one hand, it follows from (14) and (24) that, for $z\in\mathbb{N}$,
$z\int_{\mathbb{N}_{-1}}\Big{(}\log\Big{(}1+\frac{x}{z+1}\Big{)}\Big{)}^{2}\nu_{\rm br}({\rm d}x)=zd\Big{(}\log\Big{(}1+\frac{1}{z}\Big{)}\Big{)}^{2}+z\sum_{i\geq 1}\pi_{i}\Big{(}\log\Big{(}1+\frac{i}{z+1}\Big{)}\Big{)}^{2}\leq d+\sum_{i\geq 1}i\pi_{i},$ (40)
@@
On the other hand, observe that $\log(1+x)\leq\sqrt{2}x^{\alpha/2}$, $x\geq 0$. Then, it follows from the previous inequality, (14) and (${\bf H1}$) that, for $z\in\mathbb{N}$,
@@
By combining (39), (40) and (3.2), we see that there exists a constant $C_{\alpha}>0$ such that $V(z)\leq C_{\alpha}z^{2-\alpha}$, for $z\in\mathbb{N}_{0}$. Moreover, Lemma 4 (iii) implies that for any $t>0$, there exists a constant $C>0$ such that $v_{s}\leq Cs^{-1}$ for $s\in[0,t]$. Therefore, for all $s\in[0,t]$ and any $\varepsilon>0$, there exists a constant $C_{\alpha,\varepsilon,t}>0$ such that $\widehat{V}_{\varepsilon}(z_{0},s)\leq C_{\alpha,\varepsilon,t}s^{-2+\alpha}$, which clearly implies (38) (recall that $\alpha\in(1,2]$ in (${\bf H1}$)). ∎
@@
Note that the function $\widehat{V}_{\varepsilon}$ in Lemma 5 corresponds to $\hat{V}_{F,\varepsilon}$ in [3, Assumption 4.3].
Remark 1.
@@
If (${\bf H1}$) is satisfied with $\alpha=2$ (i.e., $\bm{\sigma}^{2}_{{\rm int}}<\infty$), then it follows from $\mathbf{m}_{\rm br}<\infty$, (39), (40) and (3.2) that the function $V$ in (37) is bounded on $\mathbb{N}_{0}$. In particular, the function $\widehat{V}_{\varepsilon}$ is also bounded which implies (38). Moreover, in this case, the hypothesis that $\mathbf{m}_{\rm int}<0$ is not needed and (38) is also valid under the critical cooperative condition $-c+\sum_{i\geq 1}ib_{i}=0$.
We have now all the ingredients to prove Theorem 1.
Proof of Theorem 1.

### v2

@@ 3.1 Integral representation of BPI processes
The proof is a simple adaptation of the argument used in the Proof of Theorem 5.5 in [12]. Let $Y_{t}=Z^{(z_{1})}_{t}-Z^{(z_{2})}_{t}$, for $t\geq 0$. By Proposition 3.1,
$Y_{t}=z_{1}-z_{2}+\int_{0}^{t}\int_{\mathcal{X}}(H(Z^{(z_{1})}_{s-},u,y,x)-H(Z^{(z_{2})}_{s-},u,y,x))N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x).$ (3.7)
Let $(a_{m})_{m\geq 0}$ be a decreasing sequence of positive real numbers such that $a_{0}=1$, $a_{m}>0$ for $m\in\mathbb{N}$, and $\lim_{m\rightarrow\infty}a_{m}=0$. For each $m\in\mathbb{N}$, let $g_{m}:\mathbb{R}\rightarrow\mathbb{R}_{+}$ be a non-negative continuous function with support on $(a_{m},a_{m-1})$ and satisfying $\int_{a_{m}}^{a_{m-1}}g_{m}(z){\rm d}z=1$. For each $m\in\mathbb{N}$, we define the non-negative and twice continuously differentiable function
$\phi_{m}(z)=\int_{0}^{z}\int_{0}^{y}g_{m}(x){\rm d}x{\rm d}y,\qquad z\in\mathbb{R}.$ (3.8)
@@
$G_{s}(\phi_{m},Y)\coloneqq\phi_{m}(Y_{s-}+H(Z^{(z_{1})}_{s-},u,y,x)-H(Z^{(z_{2})}_{s-},u,y,x))-\phi_{m}(Y_{s-})=0,\quad s\geq 0.$ (3.10)
Note that $Y_{0}=z_{1}-z_{2}\leq 0$ and $\phi_{m}(Y_{0})=0$. For $n\in\mathbb{N}$, define $\tau_{n}=\inf\{t\geq 0:Z^{(z_{1})}_{t}\geq n\,\,\text{or}\,\,Z^{(z_{2})}_{t}\geq n\}$. Then, by (3.7), Itô’s formula (Theorem 5.1 in Chapter 2 of [17]) and (3.10),
@@
$=\int_{0}^{t\wedge\tau_{n}}\int_{\mathcal{X}}G_{s}(\phi_{m},Y)\mathbf{1}_{\{Y_{s-}>0\}}{\rm d}sq({\rm d}u,{\rm d}y,{\rm d}x)+M_{t}^{(n,m)},$ (3.11)
where $(M_{t}^{(n,m)})_{t\geq 0}$ is martingale. Note that $|\phi_{m}^{\prime}(z)|\leq 1$, for $z\in\mathbb{R}$, which implies that $\phi_{m}$ is (globally) Lipschitz. Since we have assumed that $\sum_{i\geq 1}i\pi_{i}<\infty$ and $\sum_{i\geq 1}ib_{i}<\infty$, it follows from (3.4) and (3.6), for any $s\leq\tau_{n}$,
$\int_{\mathcal{X}}G_{s}(\phi_{m},Y)\mathbf{1}_{\{Y_{s-}>0\}}q({\rm d}u,{\rm d}y,{\rm d}x)\leq\Big{(}d+(2n-1)c+\sum_{i\geq 1}i(\pi_{i}+(2n-1)b_{i})\Big{)}Y_{s-}^{+}.$ (3.12)
Then, from (3.1), (3.12) and the dominated convergence theorem (letting $m\rightarrow\infty)$, we deduce that
$\mathbb{E}[Y_{t\wedge\tau_{n}}^{+}]\leq\Big{(}d+(2n-1)c+\sum_{i\geq 1}i(\pi_{i}+(2n-1)b_{i})\Big{)}\int_{0}^{t}\mathbb{E}[Y_{s\wedge\tau_{n}}^{+}]{\rm d}s.$ (3.13)
Then, $\mathbb{E}[Y_{t\wedge\tau_{n}}^{+}]=0$, for all $t\geq 0$. On the other hand, since $(Z^{(z_{1})}_{t})_{t\geq 0}$ and $(Z^{(z_{2})}_{t})_{t\geq 0}$ do not explode in finite time and have càdlàg paths, we deduce that $\tau_{n}\rightarrow\infty$, as $n\rightarrow\infty$. Therefore, our claim follows. ∎
Lemma 3.3.
Suppose that $\sum_{i\geq 1}i\pi_{i}<\infty$ and $\sum_{i\geq 1}ib_{i}<\infty$. Then, the BPI process is stochastically monotone, that is, for all $z_{1},z_{2}\in\mathbb{N}_{0}$ such that $z_{1}\leq z_{2}$, $\mathbb{P}(Z_{t}^{(z_{1})}\geq x)\leq\mathbb{P}(Z_{t}^{(z_{2})}\geq x)$, for all $t\geq 0$ and $x\in\mathbb{R}$.
@@
It follows from Proposition 3.2. ∎
3.2 Proof of Theorem 2.1
In this section we prove Theorem 2.1 using the framework developed in Section 4 of [3] (with $F(z)\coloneqq\log(1+z)$, for $z\in(-1,\infty)$). Recall that $\mathcal{X}\coloneqq\{1,2\}\times(\mathbb{N}\cup\Delta)\times\mathbb{N}_{-1}$. The next lemma verifies part of Assumptions 4.1 and 4.3 in [3]. We have removed the subscript $F$ from our notation and will use $\varphi$ to denote the function $b_{F}$ in [3].
Lemma 3.4.
@@
    Define the function $h(z)\coloneqq\int_{\mathcal{X}}\log\Big{(}1+\frac{H(z,u,y,x)}{z+1}\Big{)}q({\rm d}u,{\rm d}y,{\rm d}x)$, for $z\in\mathbb{N}_{0}$. Then, $h$ can be extended to $0,\infty)$ such that is locally bounded on $0,\infty)$ and locally Lipschitz on $(0,\infty)$.
@@
    Define the function $\psi(z)\coloneqq e^{-z}\varphi(e^{z}-1)$, for $z\in[0,\infty)$. Then, $\psi$ is $(0,\kappa)$ nonexpansive on $(1,\infty)$ for some $\kappa\geq 0$, i.e. for all $z_{1}>z_{2}>0$, $\psi(z_{1})\leq\psi(z_{2})+\kappa$.
@@
$\ln(1+x)\leq x.$ (3.14)
First, we prove (i). If $z=0$, then $H(0,u,y,x)=0$, for all $u=1,2$, $y\in\mathbb{N}\cup\Delta$ and $x\in\mathbb{N}_{-1}$ and thus, (i) holds. For $z\in\mathbb{N}$, (i) follows from (3.4), (3.6), (3.14) and our assumptions $\mathbf{m}_{\rm int}<0$ and $\mathbf{m}_{\rm br}<\infty$, that is,
@@
$\leq d+\sum_{i\geq 1}i\pi_{i}+zc+z\sum_{i\geq 1}ib_{i}<\infty.$ (3.15)
Next, we prove (ii). Note that by (3.4) and (3.6),
$h(z)$ $=z\int_{\mathbb{N}_{-1}}\log\Big{(}1+\frac{x}{z+1}\Big{)}\nu_{\rm br}({\rm d}x)+\frac{z(z-1)}{2}\int_{\mathbb{N}_{-1}}\log\Big{(}1+\frac{x}{z+1}\Big{)}\nu_{\rm int}({\rm d}x)=h_{\rm br}(z)+h_{\rm int}(z),$ (3.16)
@@
$h_{\rm int}(z):=-z(z-1)c\log\Big{(}1+\frac{1}{z}\Big{)}+z(z-1)\sum_{i\geq 1}b_{i}\log\Big{(}1+\frac{i}{z+1}\Big{)}.$ (3.18)
Clearly, $h$ (and thus, $h_{\rm br}$ and $h_{\rm int}$) can be extended to $0,\infty)$. In particular, (3.2) remains valid and it follows that $h$ is locally bounded on $0,\infty)$. On the other hand, note that, for $0<x\leq y$, (3.14) implies that
$|\log(x)-\log(y)|=\log\left(\frac{y}{x}\right)=\log\left(1+\left(\frac{y}{x}-1\right)\right)\leq\frac{y-x}{x}.$ (3.19)
Then, by (3.19) and our assumptions that $\mathbf{m}_{\rm int}<0$ and $\mathbf{m}_{\rm br}<\infty$, it is not difficult to deduce that $h$ is locally Lipschitz on $(0,\infty)$, which concludes with the proof of (ii).
We now prove (iii). By (3.14) and since $\mathbf{m}_{\rm br}<\infty$, $h_{\rm br}$ is a bounded function on $0,\infty)$. On the other hand, by (3.14), for $z\in[0,\infty)$, $h_{\rm int}(z)\leq(z-1)\mathbf{m}_{\rm int}$. In particular, since $\mathbf{m}_{\rm int}<0$, $h_{\rm int}(z)\rightarrow$-$\infty$, as $z\rightarrow\infty$, which implies (iii).
@@
$\left|\frac{1}{x}-\frac{1}{y}\right|=\frac{|x-y|}{xy}.$ (3.20)
Then, by the triangle inequality and (3.20), we have that, for $z\in(1,\infty)$,
$\Big{|}-z(z-1)c\log(1+1/z)+cz\Big{|}$ $\leq z(z-1)\Big{|}c\int_{0}^{-1}\Big{(}\frac{1}{z+1+y}-\frac{1}{z-1}\Big{)}{\rm d}y\Big{|}\leq 2c.$ (3.21)
[…]
Let $z_{0}>1$ be such that (iii) in Lemma 3.4 is satisfied. Recall the definition of the function $\varphi$ in Lemma 3.4 (iii). By Lemma 3.4 (i)-(ii), $\varphi$ is well-defined and locally Lipschitz on $D\coloneqq(z_{0},\infty)$. We introduce the flow $\phi$ associated to $\varphi$ and defined for $z\in D$ as the unique solution of
$\quad\frac{\partial}{\partial t}\phi(z,t)=\varphi(\phi(z,t)),\qquad\textrm{with}\quad\phi(z,0)=z,$ (3.22)
@@
$v_{t}\coloneqq\phi(\infty,t)=\lim_{z\rightarrow\infty}\phi(z,t),$ (3.23)
for any $t\in[0,T(\infty))$. By Lemma 3.4 (iii), for $z\in D$, $\varphi(z)<0$ and for any $t<T(z)$,
$\int_{z}^{\phi(z,t)}\frac{{\rm d}x}{\varphi(x)}=t.$ (3.24)
Lemma 3.5.
@@
$-\int_{z}^{\infty}\frac{{\rm d}x}{\varphi(x)}<\infty.$ (3.25)
@@
The first claim follows from the definition of $\varphi$, (3.16) and the dominated convergence theorem. The second claim follows clearly from the first claim. ∎
By Lemma 3.5, we have that $v_{t}$ satisfies
$-\int_{v_{t}}^{\infty}\frac{{\rm d}x}{\varphi(x)}=t,$ (3.26)
@@
$v_{t}=\inf\Big{\{}z\in D:\int_{z}^{\infty}\frac{{\rm d}x}{-\varphi(x)}<t\Big{\}}<\infty,$ (3.27)
for any $t\in(0,T(\infty))$. Following Section4 in [3], one says that the dynamical system instantaneously comes down from infinity. The function $t\in[0,T(\infty))\rightarrow v_{t}\in\overline{\mathbb{R}}$ is continuous when $\overline{\mathbb{R}}$ is endowed with the distance $\overline{d}(x,y)=|e^{-x}-e^{-y}|$, for $x,y\in\overline{\mathbb{R}}$.
Lemma 3.6.
@@
By (3.23), $v_{t}\in D$ and thus $v_{t}>0$, for $t\in[0,T(\infty))$. By Lemma 3.4 (iii), $\varphi(\cdot)$ is strictly positive on $D$. Then, Lemma 3.5 implies that $z\mapsto g(z)=-\int_{z}^{\infty}\frac{{\rm d}x}{-\varphi(x)}$ maps $D=(z_{0},\infty)$ bijectively to $(0,-\int_{z_{0}}^{\infty}\frac{{\rm d}x}{-\varphi(x)})$. Moreover, $g$ is decreasing on $(z_{0},\infty)$. Then, it should be clear that $\lim_{t\downarrow 0}v_{t}=\infty$, which proves (i). (ii) follows by (3.26) and the fundamental theorem of calculus; recall that for $z\in D$, $\varphi(z)<0$. From (3.26), we see that $tv_{t}=v_{t}\int_{v_{t}}^{\infty}\frac{{\rm d}x}{-\varphi(x)}$, for $t\in(0,T(\infty))$. Then, L’Hôpital’s rule and Lemma 3.5 imply (iii), that is,
$\lim_{t\downarrow 0}-\mathbf{m}_{\rm int}tv_{t}=-\mathbf{m}_{\rm int}\lim_{z\rightarrow\infty}z\int_{z}^{\infty}\frac{{\rm d}x}{-\varphi(x)}=-\mathbf{m}_{\rm int}\lim_{z\rightarrow\infty}\frac{\frac{1}{\varphi(z)}}{-\frac{1}{z^{2}}}=1.$ (3.28)
@@
The following lemma shows that condition (21) in Assumption 4.3 of [3] is satisfied.
Lemma 3.7.
@@
$V(z)\coloneqq\int_{\mathcal{X}}\Big{(}\log\Big{(}1+\frac{H(z,u,y,x)}{z+1}\Big{)}\Big{)}^{2}q({\rm d}u,{\rm d}y,{\rm d}x),$ (3.29)
@@
$\int_{0}^{t}\widehat{V}_{\varepsilon}(z_{0},s){\rm d}s<\infty,$ (3.30)
@@
$V(z)=z\int_{\mathbb{N}_{-1}}\Big{(}\log\Big{(}1+\frac{x}{z+1}\Big{)}\Big{)}^{2}\nu_{\rm br}({\rm d}x)+\frac{z(z-1)}{2}\int_{\mathbb{N}_{-1}}\Big{(}\log\Big{(}1+\frac{x}{z+1}\Big{)}\Big{)}^{2}\nu_{\rm int}({\rm d}x).$ (3.31)
On the one hand, observe that $\log(1+x)\leq\sqrt{2x}$, $x\geq 0$. Then, it follows from (3.1) and (3.14) that, for $z\in(0,\infty)$,
$z\int_{\mathbb{N}_{-1}}\Big{(}\log\Big{(}1+\frac{x}{z+1}\Big{)}\Big{)}^{2}\nu_{\rm br}({\rm d}x)$ $=zd\Big{(}\log\Big{(}1+\frac{1}{z}\Big{)}\Big{)}^{2}+z\sum_{i\geq 1}\pi_{i}\Big{(}\log\Big{(}1+\frac{i}{z+1}\Big{)}\Big{)}^{2}$
$\leq 2\Big{(}d+\sum_{i\geq 1}i\pi_{i}\Big{)},$ (3.32)
@@
On the other hand, observe that $\log(1+x)\leq\sqrt{2}x^{\alpha/2}$, $x\geq 0$. Then, it follows from the previous inequality, (3.1) and ($\bf H1$}) that, for $z\in(0,\infty)$,
@@
By combining (3.31), (3.2) and (3.33), we see that there exists a constant $C_{\alpha}>0$ such that $V(z)\leq C_{\alpha}(z^{2-\alpha}+1)$, for $z\in[0,\infty)$. Moreover, Lemma 3.6 (iii) implies that for any $t>0$, there exists a constant $C_{t}>0$ such that $v_{s}\leq C_{t}s^{-1}$ for $s\in[0,t]$. Therefore, for all $s\in[0,t]$ and any $\varepsilon>0$, there exists a constant $C_{\alpha,\varepsilon,t}>0$ such that $\widehat{V}_{\varepsilon}(z_{0},s)\leq C_{\alpha,\varepsilon,t}(s^{-2+\alpha}+1)$, which clearly implies (3.30) (recall that $\alpha\in(1,2]$ in ($\bf H1$})). ∎
@@
Note that the function $\widehat{V}_{\varepsilon}$ in Lemma 3.7 corresponds to $\hat{V}_{F,\varepsilon}$ in Assumption 4.3 of [3].
Remark 3.8.
@@
If ($\bf H1$}) holds with $\alpha=2$, then $\bm{\sigma}^{2}_{{\rm int}}<\infty$. Since we have already assumed that $\mathbf{m}_{\rm br}<\infty$, it follows from (3.31), (3.2) and (3.33) (with $\alpha=2$) that the function $V$ in (3.29) is bounded. In particular, the function $\widehat{V}_{\varepsilon}$ is also bounded, which implies (3.30).
Furthermore, the proof shows that the result of Lemma 3.7 holds true even without the condition that $\mathbf{m}_{\rm int}<0$, as long as ($\bf H1$}) is satisfied.
We have now all the ingredients to prove Theorem 2.1.
Proof of Theorem 2.1.